OpenStudy (anonymous):

PLEASE HELP ME---A wheelchair access ramp has an angle of elevation of 24°. If the ramp reaches to the top of a 30 inch high porch, how long is the ramp?

OpenStudy (debbieg):



OpenStudy (debbieg):

You know an angle and the opposite side, you want the adjacent side. Which trig function relates those 3 parts?

OpenStudy (debbieg):

Set up the equation, then solve it for L. :)

OpenStudy (debbieg):

Whoh, wait, I drew that wrong... hang on....

OpenStudy (debbieg):



OpenStudy (debbieg):

You know an angle and the opposite side, you want the hypotenuse. Which trig function relates those 3 parts? Proceed as outlined above... :)

OpenStudy (anonymous):

um it would be sine right? sin24?

OpenStudy (debbieg):

Well, that's the right function, yes. Now what's the equation?

OpenStudy (debbieg):

sin24*={what?}/{what?}

OpenStudy (anonymous):

sin24= 30/ x???

OpenStudy (debbieg):

Yup... exactly! Now solve that puppy for x.

OpenStudy (anonymous):

i thnk i did something wrong in the solving because i got a negative number

OpenStudy (debbieg):

I think your calculator is in radian mode.

OpenStudy (debbieg):

Because I'm kinda psychic like that. :)

OpenStudy (anonymous):

oh hahahah i think you're right lol

OpenStudy (anonymous):

ok now its on deg but i got a really small number

OpenStudy (debbieg):

hmmm.... can you show me exactly what you entered in your calc?

OpenStudy (anonymous):

i put in sin 24 and i got 0.40673 then i divided it by 30

OpenStudy (anonymous):

i did it wrong huh?

OpenStudy (debbieg):

Yup... hold on, was having technical problems...

OpenStudy (anonymous):

ok haha

OpenStudy (debbieg):

You have: \(\Large \sin24^*=\dfrac{30}{x}\) To solve that for x, you need to MULTIPLY both sides by x then DIVIDE both sides by sin(24*)

OpenStudy (debbieg):

But YOU ended up with sin(24*)/30.... that's backwards.

OpenStudy (debbieg):

That would be giving you 1/x. :) \(\Large \dfrac{1}{30}\cdot\sin24^*=\dfrac{30}{x}\cdot\dfrac{1}{30}\) \(\Large \dfrac{\sin24^*}{30}=\dfrac{1}{x}\) See what I mean?

OpenStudy (anonymous):

ohh yeah i got you

OpenStudy (debbieg):

Here is what you MEANT to do: \[\Large \dfrac{x}{\sin24^*}\cdot\sin24^*=\dfrac{30}{x}\cdot\dfrac{x}{\sin24^*}\] \(\Large x=\dfrac{30}{\sin24^*}\)

OpenStudy (anonymous):

ohh i get it

OpenStudy (anonymous):

so the answer would be 73.75

OpenStudy (debbieg):

That's what I got. :)

OpenStudy (anonymous):

omg yay haha thanks so much!

OpenStudy (debbieg):

You're welcome, happy to help! :)